OpenStudy (anonymous):

Rewrite the equation of a line in point-slope form to slope intercept form: y + 2 = –2(x – 5)

OpenStudy (anonymous):

are you still there?

OpenStudy (anonymous):

Well slope intercept form is y=mx+b So looking at your equation: Y+2=-2(x-5) You wold multiply everything in the parenthesis by -2 to get: Y+2=-2x+10 You would then isolate y by subtracting two from both sides to give you your answer of: Y=-2x+8

OpenStudy (anonymous):

y=-2x+8 please remember to give medals ^.^